OpenStudy (anonymous):

Evaluate (ab)2 for a = 4 and b = 3. A. 36 B. 24 C. 81 D. 144

OpenStudy (anonymous):

a*b=?

OpenStudy (anonymous):

You plug in what a and b are into the parentheses. Multiply it. Then multiply the 2 and then you get your answer.

OpenStudy (anonymous):

4*3=12 12^2=?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

no 12*12=?

OpenStudy (anonymous):

it's 12 *2

OpenStudy (anonymous):

\[(ab)^2 \] or (ab)*2

OpenStudy (anonymous):

@jordyn77 if it was times there would been no reason to but the parentheses thats what i thought

OpenStudy (anonymous):

Yeah we need to be sure though, because if it is ^2 then you're right (:

OpenStudy (anonymous):

@jennyjewell25 can you check what i asked

OpenStudy (anonymous):

24